OpenStudy (anonymous):

Compute Δy and dy for the given values of x and dx = Δx. (Round your answers to three decimal places.) Compute Δy and dy for the given values of x and dx = Δx. (Round your answers to three decimal places.) y = radical(x) x = 1, Δx = 1 find Δy and dy

OpenStudy (anonymous):

I don't know how to find Δy

OpenStudy (anonymous):

dy = 1/2 but I don't know how to find delta y

ganeshie8 (ganeshie8):

\[\Delta y = y_2 - y_1 \]

ganeshie8 (ganeshie8):

x starts at 1, so \(\large x_1 = 1\) \(\large x_2 = x_1 + \Delta x = 1 + 1 = 2\)

ganeshie8 (ganeshie8):

use them to evaluate \(\large y_2 \) and \(\large y_1\)